A 3.0-kg cylinder falls vertically from rest in a
Answer #1

use,


m*g*h=1/2*m*v^2


==> v=sqrt(2*g*h)


v=sqrt(2*9.8*6)


v=10.84 m/sec or 11 m/sec <-----